OpenStudy (anonymous):

If each angle in a pentagon is the same measure, what is that measure? 36° 540° 180° 108°

OpenStudy (anonymous):

108

OpenStudy (lgbasallote):

i suggest you use the formula 180(n-2) where n is the number of sides

OpenStudy (anonymous):

Externals always add to 360 (convex)

OpenStudy (lgbasallote):

oh yeah then divide the product by n

OpenStudy (anonymous):

You can figure that out by doing this: 5-2=3 180*3= 540 540/5=108

OpenStudy (anonymous):

180-360/5

OpenStudy (anonymous):

pentagon= 5 sided , 5(x)=360 5x=360 x=360/5 the degree measures of the sides is 360/5